# EventBind: Learning a Unified Representation to Bind Them All for Event-based Open-world Understanding

### v4

@@ 4.4 Extension Experiment: Event Retrieval
Figure 6: Visual examples for the event retrieval task.
EventBind can be expanded to image-to-event and text-to-event retrieval tasks. See more details about event retrieval and results discussions in Suppl. We initially perform the event retrieval task by employing both text and image queries. To evaluate performance, we utilize Recall@1, Recall@5, and Recall@10 metrics commonly employed in retrieval tasks [60, 11]. In Fig. 6, our EventBind shows remarkable retrieval performance (99.01% Recall@1 for text query and 96.70% Recall@1 for image query) with only 20 shot training examples, demonstrating its remarkable capabilities in few-shot learning. Notably, our model excels in text-to-event and image-to-event retrieval, achieving recall rates close to 100.00% on Recall@1 after fine-tuning. This exceptional performance demonstrates that EventBind effectively establishes a unified representation space with aligned event, images and text embeddings. Besides, In Fig. 6, we present the Top-5 event data retrieved from the N-Caltech101 dataset using randomly selected images or class texts as queries. All the retrieved event data exhibit a very high degree of similarity to the input text query, showcasing the effectiveness of our proposed EventBind framework.
5 Conclusion and Future Work

### v6

@@ C. Additional Experiment Result
Event Retrieval Numerical Results We utilize Recall@1, Recall@5, and Recall@10 metrics commonly employed in retrieval tasks [69, 11]. In Fig. 7, our EventBind shows remarkable retrieval performance (99.01% Recall@1 for text query and 96.70% Recall@1 for image query) with only 20 shot training examples, demonstrating its remarkable capabilities in few-shot learning. Notably, our model excels in text-to-event and image-to-event retrieval, achieving recall rates close to 100.00% on Recall@1 after fine-tuning. This exceptional performance demonstrates that EventBind effectively establishes a unified representation space with aligned event, images and text embeddings. To ease future comparison, we report all those numbers in Tab. 16.